OpenStudy (anonymous):

Create an equation. Use the graph below to create the equation of the rainbow parabola.

OpenStudy (anonymous):

it's an upside down parabola

OpenStudy (anonymous):

@Loser66 @Luigi0210

OpenStudy (loser66):

what is the vertex-form of a parabola?

OpenStudy (anonymous):

um... (h,k) is the vertex

OpenStudy (anonymous):

y=a(x-h)^2-k @Loser66 ??

OpenStudy (loser66):

yes, but +k, not -k and your vertex is ???

OpenStudy (anonymous):

(0,36)

OpenStudy (loser66):

ok, so, it is ?? y = ??

OpenStudy (loser66):

replace h, k in

OpenStudy (anonymous):

um...y=x^2+36?

OpenStudy (loser66):

where is your a??

OpenStudy (loser66):

y =a x^2 +36 , ok??

OpenStudy (anonymous):

kk!

OpenStudy (loser66):

Now, pick a point on the graph. (either of them, not vertex)

OpenStudy (anonymous):

um -6?

OpenStudy (loser66):

a point means (.., ...) not just -6

OpenStudy (anonymous):

(0,6)

OpenStudy (loser66):

you don't have that point!! , you have (-6,0) or (6,0)

OpenStudy (loser66):

which one do you choose??

OpenStudy (anonymous):

whoops! um...(6,0) :)

OpenStudy (loser66):

ok, that is x =6 , y =0 right? plug that point into y = ax^2 +36, what do you have?

OpenStudy (triciaal):

an upside down parabola / like a rainbow means the value of a is negative

OpenStudy (anonymous):

y=-x^2+36

OpenStudy (loser66):

@triciaal please, be patient, it will be there!!

OpenStudy (loser66):